OpenStudy (anonymous):

in a certain year the amount A of garbage in pounds produced after t days by an average household is given by A=8.5t

OpenStudy (anonymous):

Graph the equation for t>0. How many days did it take to produce 200 pounds of garbage?

OpenStudy (ghazi):

to produce 200 pounds of grabage ...days required= 200/8.5

OpenStudy (anonymous):

that would equal out to be 23.5 days right?

OpenStudy (ghazi):

right

OpenStudy (anonymous):

Okay so how would I graph it?

OpenStudy (ghazi):

to plot graph you've to use different values of t and then find values of A...likt at t=1 A= 8.5....similarly substitute the different values of t and find A then plot it on graph....

OpenStudy (ghazi):

 for different t find A and plot it

OpenStudy (anonymous):

Okay, you kind of lost me there. So I would have to substitute 1 from 8.5?

OpenStudy (ghazi):

no...take t= 1..put it in A=8.5*t=8.5*1=8.5...then take t= 1/2...therefore A= 8.5*1/2=4.25...and so on :)

OpenStudy (anonymous):

oh okay, I see now

OpenStudy (ghazi):

can you do it now? and try to use smaller value of t ..it would be easier to plot

OpenStudy (anonymous):

I think I got it

OpenStudy (anonymous):

Okay I don't have it, I got 4.25 but then I don't know what to do after that.

OpenStudy (ghazi):

well you've to plot it on graph , now

OpenStudy (anonymous):

Oh, wow. Just shows how much I know about math

OpenStudy (ghazi):

 it will come like this...but it's just a rough sketch